OpenStudy (anonymous):

I need help solving this equation. 4-8x/x+1 = 8/x+1 can anyone help me?

OpenStudy (amistre64):

is it: (4-8x)/(x+1) = 8/(x+1) ?

OpenStudy (amistre64):

or.. (4-8x)/(x+1) = (8/x)+1?

OpenStudy (amistre64):

first ones easy: denominators dont count.. much. 4-8x = 8 -8x = 8-4 -8x = 4 x = (-1/2)

OpenStudy (nowhereman):

\[4 - \frac{8x}{x} + 1 = \frac{8}{x} + 1 \] \[ ⇔ x \neq 0 ∧ -3 = \frac{8}{x} + 1 \] \[ ⇔x \neq 0 ∧ x = \frac{8}{-4} ⇔ x = -2\]

OpenStudy (anonymous):

The problem is 4- 8x over x+1 = 8 over x+1. I have to solve it.

OpenStudy (amistre64):

i solved it fer ya; x = -(1/2) or 1.5 if you want a decimal

OpenStudy (amistre64):

that 1 was supposed to be a (-) lol.... a aint a typist

OpenStudy (anonymous):

thank you so much I'm not much of a mathamatian. sorry if spelled wrong.LOL

OpenStudy (amistre64):

i believe they are called "mathmatickers" but I could be mistaken :)